Please help me
"Amie is five years younger than his brother Benny. In four years from now, Benny will be twice as old as Amie. Find their present ages"

Answer by Edwin McCravy(20060)


"Amie is five years younger than his brother Benny.
Amie is Benny's age minus 5.

Amie = (B-5)

In four years from now, Benny will be twice as old as Amie.
Four years from now Benny will be B+4 and Amie will be (B-5)+4.

Then B+4 will equal 2[(B-5)+4]

     B+4 = 2[(B-5)+4]

Solve that for Benny's age now.

Then substitute what you get in

Amie = (B-5)

to find Amie's age now.

Edwin